OpenStudy (anonymous):

(1/2)x + (1/4)x = 16

OpenStudy (anonymous):

First do + because we have only x s : Means : (1/2+1/4)x

OpenStudy (anonymous):

A.16 B.21.3 C.32 D.64

OpenStudy (anonymous):

these are the choices

OpenStudy (anonymous):

divide each side by 2

OpenStudy (anonymous):

Okay, then what

OpenStudy (anonymous):

divide each side by 4

OpenStudy (anonymous):

what would the answer be

OpenStudy (anonymous):

u tell me?

OpenStudy (anonymous):

Im working on it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

This is a little tough, i havent done this in a while

OpenStudy (anonymous):

im sorrry i meant multiply everything on the left side by 2.

OpenStudy (anonymous):

and the right side

OpenStudy (anonymous):

ohhhh, well that makes this easier,

OpenStudy (anonymous):

i got 4/4

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

no... okay what did you get when you multiplied everything by 2

OpenStudy (anonymous):

hold on, 2/4+ 2/8

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

i said everything by 2/1 or 2 so u should get... x+ 2/4x=32

OpenStudy (anonymous):

2(1/2x) + 2(1/4x) = 2(16)

OpenStudy (anonymous):

2/2x + 2/8x = 32

OpenStudy (anonymous):

correct! now do the rest and see what u get.

OpenStudy (anonymous):

4/10x=32

OpenStudy (anonymous):

actually no its 2/2x + 2/4x = 32

OpenStudy (anonymous):

2/4x not 2/8x

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

u tell me... if you have 2/2x + 1/4 x = 16 do the rest. its pretty simple.

OpenStudy (anonymous):

3/6x=16

OpenStudy (anonymous):

*i mean if you have 2/2x + 2/4 x = 16

OpenStudy (anonymous):

so 2/2 is also 1... so change 2/2x to 1x or x

OpenStudy (anonymous):

then what

OpenStudy (anonymous):

now you would have x + 2/4x = 32

OpenStudy (anonymous):

WE have : (1.2+1.4)x=16 2 .6x=16 x= 16/26/10 X = 160/26

OpenStudy (anonymous):

so,which answer is it(answer choices at the top)

OpenStudy (anonymous):

No one ! I m sure for my answer . I think you cant say the question well . Or the question is fall !:

OpenStudy (anonymous):

C